OpenStudy (rossmeador):

Find the explicit formula a17=-145 and a34=-315

OpenStudy (unklerhaukus):

\(a_{17}=-145\) and \(a_{34}=-315\) these points are \(34-17=\cdots\) far apart in the series and \(-315-(-145)=\cdots\) far apart in value the relationship between these two differences is easy to see

OpenStudy (rossmeador):

-10, so how would I find the explicit formula?

OpenStudy (rossmeador):

an=a1 times -10?

OpenStudy (unklerhaukus):

not quite

OpenStudy (rossmeador):

Can you repost that one more time please? It came up for 5 seconds but for some reason closed.

OpenStudy (unklerhaukus):

yeah i made a mistake so i scrapped it

OpenStudy (rossmeador):

Ok

OpenStudy (unklerhaukus):

the points are 17 apart in the series and -170 apart in value so the relationship between element in series : value 17 : -170 \[\qquad\qquad\frac{17}{-170}=\frac1{10}\]

OpenStudy (unklerhaukus):

* \( -\frac 1{10}\)

OpenStudy (rossmeador):

Ahhh. So would it be an=a1 times -1/10?

OpenStudy (unklerhaukus):

how to get the 5?

OpenStudy (anonymous):

d=(a34-a17)/(34-17)=-10; an=a1+(-10)*(n-1); a17-a1=16*d;

OpenStudy (rossmeador):

What do you mean by the 5?

OpenStudy (unklerhaukus):

if we are taking away 10 every time we move along in the sequence how can we get values like \(-145,-315\)

OpenStudy (unklerhaukus):

value:element (n) \[-10:1\]is\[-10\]

OpenStudy (anonymous):

I just solved it.

OpenStudy (rossmeador):

That's right!

OpenStudy (unklerhaukus):

i don't see how you got your second line @Farmer_Wu and i dont understand your third, can you explain a little bit ?

OpenStudy (anonymous):

It is just a formula you can just remmenber it. the third is just use the same way of the first.